eve shift 7 Jan 00


Subject: eve shift 7 Jan 00
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:09:48 CST


                Evening Shift 7 Jan 00
 
On Shift: Sydney,Marj

Hours of beam available at all intensities: 7.0
Hours of beam available at nominal intensity: 7.0
Hours of E799 data (special runs not included): 6.0
Hours of special runs (laser scans, muon runs...): 0.1
Accounting of lost time:
  KFRENDs, Pipeline reset, starting and stopping runs

========================================================================
Tapes brought to FCC this shift: None
  

========================================================================
Accesses during shift: None

========================================================================
Runs ended this shift:

15427 E799 data 19M events
15430 Muon run, 100K events
15431 E799 data, 1.1M events but DC1 T0's are bad
15432 E799 data 2.6M events
15433 E799 data, ongoing
 
=======================================================================
Overview of shift: Fairly smooth E799 data. A rash of lost permits
around 21:00, the problem went away with multiple KFRENDs. Also took
a muon run (with BA HV's as per Mike Arenton), changed NM4AN polarity
to positive. The new T0's loaded around 18:00 had a problem, went back
to old file.

   Rick K put quartz bar into BA33 channel.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

15427 Run ongoing as shift starts.

15427 370,371 KTEVANA errors, KFREND

        372 pipeline errors, reset pipline several times

        380 resume data-taking after pipeline reset

15427 394-400 17:00 beam off for several spills, then returns

15427 405-406 17:15 Ktevana errors again, KFREND
         
15427 17:35 Beam off, linac klystron problems, no est.
                        End run, load new tapes, prepare for muon run,
                        NM4AN changed to positive.

                18:00 Load new T0's while beam is off.

15430 19:00 beam returns, muon run, 3 spills

15431 19:20 Beam 799 data!

                19:30 Most displays are hosed, run Fix_display
                        Need to drt_fresh_start at next start of run

15431 20:00 End 15431, DC1 had very low efficiency. Traced to
                        bad T0's in the new file, reload T0's from run 15268

15432 20:10 Start new run, now DC eff's look OK.

15432 64-74 21:25 Trigger hangs, multiple KFREND's, power
                        cycles, disable triggers, re-enable, then
                        triggers work. Continue taking data.

15432 21:45 Rick K thinks he smells smoke, but we find
                        nothing. It did appear, however, that ktev2 had
                        stopped processing. End run and drt_fresh_start.
 
15433 22:00 Start new run, all seems OK.
 
                22:10 Rick K discovers that the H-bank was being read out,
                        he fixes it so that at next KFREND it will not
                        be read out.

                        Rick connects quartz bar PMT to channel
                        33 of the BA. Need to click "reset quartz" button
                        each time you do KFREND or start a new run.
                
========================================================================
Items to Follow up on Next Shift
 
  Keep an eye on DCOR LV power supply for DC2, it is reading low (4.8V).
  
  Label 30 QKE tapes for short reg run and other special runs
 
  Don't forget to hit "reset quartz" button after each KFREND.



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:09:52 CST